Calculate the volume change (in Liters) during the combustion of propane in which ΔE = -3083.67 kJ and ΔH = -3087.12 kJ at a constant pressure of exactly one atm and constant temperature.

C3H8(g) + 5 O2(g) ↔ 3 CO2(g) + 4 H2O(l)


The product of pressure and volume change should give units of L*atm. It can be shown that 1 L*atm = 101 J.
Answer #1

By first law of the mo alynamics DE= qtw - DE 9-POW At constant pressure q = Dn. Thus, IDE= Ah-PAV, Given, DE2 - 3083.67 kJ =If you have any query regarding this you may ask in comments